Good day,please help me please.
where R=8.314 L⋅kPa/(K⋅mol)In Kp=Kc(RT)Δn
For the reaction
X(g)+3Y(g)⇌3Z(g)
Kp = 1.24×10−5 at a temperature of 169 ∘C
Calculate the value of Kc.I calculated to be 4.49x10^-4 but says incorrect.Please help
Thank you very much.
1
Expert's answer
2020-08-12T08:57:59-0400

Solution.

"Kp = Kc \times (RT)^{\Delta n}"

"\Delta n = 3 - (1+3) = -1"

"T(K) = T(^oC) + 273"

T(K) = 442 K

"Kc = \frac{Kp}{(RT)^{\Delta n}}"

"Kc = 4.56 \times 10^{-2}"

Answer:

"Kc = 4.56 \times 10^{-2}"